OpenStudy (gandalf123):

Have you ever seen rational exponents? Rational exponents are designed to fit within the properties of exponents. That is, (31/2)2 is equal to 31, or 3. What operation does the power 1/2 represent? Can you think of operations that other fractional powers represent, such as 1/3, 1/4 and 2/3?

jhonyy9 (jhonyy9):

the exponent 1/2 mean squarroot of ... for example x^(1/2) = sqrt(x) x^(1/3) = cuberoot(x)

jhonyy9 (jhonyy9):

or exponent 2/3 of x x^(2/3) = cuberoot(x^2)